OpenStudy (anonymous):

Which of the following is equivalent to the proportion 2 over 3 equals 3 over y? A. 3y = (2)(3) B. 2y = (2)(3) C. 2y = (3)(3) D. 3y = (3)(3)

OpenStudy (lgbasallote):

hint: \[\frac 23 = \frac 3y\] solve for y does that help?

OpenStudy (anonymous):

somewhat , just need my mind refeshed on how to find y .

OpenStudy (lgbasallote):

first...cross multiply...do you remember how to do that?

OpenStudy (anonymous):

3 x 3 ? 2 x y ?

OpenStudy (lgbasallote):

right. 2*y = 3*3 do you know how to simplify this equation?

OpenStudy (anonymous):

no /:

OpenStudy (lgbasallote):

nevermind...check your choices

OpenStudy (anonymous):

so my answer would be C.

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

thanks !!

OpenStudy (lgbasallote):

welcome

OpenStudy (anonymous):

C is the answer and cross multiply 3*3 = 2*y